# AM-RADIO: Agglomerative Model – Reduce All Domains Into One If you can’t beat them, join them (and then beat them)

### v3

@@ 3.3.1 Loss Balancing
Due to the number of possible combinations of loss weights between the different teachers, and even which teachers, and possible formulations of loss functions, we mostly opted toward naive loss balancing with all teachers equally weighted for spatial features. For summary features, we have $\lambda_{CLIP}=\lambda_{DINO}=1$ and $\lambda_{SAM}=0.1$ owing largely to SAM’s low performance on k-NN classification. We then also set $\gamma_{i}=1$ by default.
We did experiment with automatic loss balancing using predicted uncertainty [10], AdaLoss [24] (momentum 0.99) and separately with AMTML-KD [37], as ways to learn the balance of $\lambda_{i}$ and $\gamma_{i}$. In the case of AMTML-KD, the model would always collapse its entire weight around the CLIP teacher and would yield worse results than naive manual balancing. Based on the results in table 3, there is very little advantage to the more exotic balancing schemes, so we opt for the ”Naive” method throughout the rest of the paper.
@@
Table 3: Loss term balancing methods comparison. We use a ViT-B/14 student, and CLIP+DINOv2 teachers. We found that AdaLoss produces the best results on the ImageNet tasks, but the worst on ADE20K. Ultimately, they’re all really close.
@@ 4 Implementation Details
Performing multi-teacher distillation is not trivial due to a mismatch in feature dimensions, different input resolution, different concepts for loss computation, challenges in fitting multiple teachers into a single GPU. We address these issues next.
General. We train all student models using the AdamW [40] optimizer, batch size 1024, and cosine annealing learning rate schedule and base learning rate of $31e-$ 3. We train for 600k steps, resulting in 614M total examples seen, approximately 2% of the number of examples seen by OpenCLIP models for training from scratch. We get our CLIP teacher models from the OpenCLIP model zoo [^1] https://github.com/mlfoundations/open_clip, our DINO-V2 models from their model zoo  [^2] https://github.com/facebookresearch/dinov2, and SAM from their model zoo  [^3] https://github.com/facebookresearch/segment-anything. For our best student models, we train using the MetaCLIP [59] ViT-H/14 OpenCLIP model, the ViT-g/14 DINO-V2 model, and the Huge SAM model. We apply random scale + cropping  [^4] https://timm.fast.ai/random_resized_crop to both student and teacher inputs (e.g. all models see the same image). We chose the DataComp-1B dataset due to it having the highest quality results of the web-scale datasets we had access to.
@@
  • Standard ViT [14] architecture to match the architecture of teachers.
  • Efficient architecture variants prioritize high throughput on GPUs. See Section 5.1.
Multi-scale Teachers. Both CLIP and DINO primarily operate at a resolution of $224^{2}$, which makes it trivial to employ both to train a student model at the same resolution. SAM, however, operates at an input resolution of $1024^{2}$, which forces us to employ a few techniques.
Rank/Teacher Partitioning. We group teacher models by (batch_size, resolution), and then distribute the groups to different distributed ranks, such that each GPU processes a consistent batch size and input resolution. We also sample groups at different rates. For our training setups that include SAM, we train with 64 GPUs (ranks), half of which get the CLIP+DINO group with batch size 32 per GPU and input resolution 224, and the other half get SAM with batch size 4 per GPU and input resolution 1024. This results in an effective batch size of 1,152. For CLIP+DINO training, we use 32 GPUs, resulting in batch size 1024.
Multi-Resolution ViTs. Many of our student models use ViT [14] as the base vision architecture. Traditionally, ViTs use a learned position embedding for each input patch in an image, which in turn enforces that the model always operate at a constant resolution. We employ the Cropped Position Embedding (CPE) [29] augmentation with the number of positions being equal to $64^{2}$ or the size of the teacher, whichever is larger. The position embeddings are then randomly cropped and interpolated to match the number of input patches for the student model. Even when training with CLIP+DINO at 224 resolution, we found that this technique results in a negligible drop (Table 4) in holistic metrics, but improved semantic segmentation linear probing mIOU. For heterogeneous-resolution teachers, this is a seamless technique that allows ViT to operate at arbitrary resolutions within some envelope.
@@
OpenCLIP H/14 81.10 40.04
DINO-v2 L/14 83.44 46.55
RADIO ViT-L/14 Non-CPE 82.96 47.30
RADIO ViT-L/14 CPE 82.84 48.52
Table 4: Comparing identical RADIO-ViT models, with and without the CPE [29] formulation, distilled from DINO-v2 ViT L/14 and OpenCLIP ViT H/14 teachers. For ADE20K segmentation, we use a simple linear head setup with the image processed at $518^{2}$ resolution. Our models were trained using the LAION-400M dataset for this experiment. Notably, the student model only ever trains at $224^{2}$ resolution, but is able to generalize to $518^{2}$ and even outperform its DINO-v2 teacher on semantic segmentation.
High-Resolution ViT Student. In SAM, they employ the ViTDet [33] architecture as a way to reduce the computational and memory burden of ViT models at high-resolution. A key part of the architecture is to convert the MHSA layers to windowed attention, and then have a normal global attention layer every $k$ layers. It turns out that this architecture choice can easily be adapted into a training augmentation, where we sample a window size from a set of possible window sizes, and for a given step, the windowed layers have that size. This allows us to meaningfully reduce the computational burden of training the student model with the SAM teacher, and, as we make the window size flexible, it provides an additional scaling mechanism during inference where the same model may be used at different degrees of computation. Table 8 shows the resulting quality of the same model being run at different window sizes.
Student/Teacher Downsampling Mismatch. When the student and teacher downsample images through their processing stack at different rates, it results in the output feature vectors having different resolutions. For example, if the teachers use a ViT-H/14 architecture and student a ViT-H/16, it means that the student outputs a $14^{2}$ feature map, and the teachers a $16^{2}$ feature map. For $L_{\text{features}}$ we bilinearly interpolate the student outputs to match the resolution of the teacher features in order to apply the loss to each output vector.
Summary Loss Feature Zero Shot k-NN ADE20K VOC GQA SQA TextVQA VQAv2
CLS token 0.07 0.09 78.55 83.91 49.01 83.51 70.46 65.89 53.36 77.66
Avgpool 0.05 0.12 80.12 83.83 38.36 77.04 70.92 65.84 54.26 78.28
Table 5: Comparing identical ViT models, with CLS token and average pooling summarization. Summary Loss - training summary loss against CLIP teacher, Feature - training feature loss against DINOv2 teacher.
Feature Summarization. In 3.3 we explained how teacher summary features are extracted using the “class token” of their respective ViT models. We now turn our attention to the summarization of student features. ViTs have 2 options: (i) a separate summarization “CLS” token or (ii) average pooling patch tokens. We evaluate both options in Table 5. We observe that average pooling improves summary loss, but has a more significant detrimental effect on the feature loss. Given the importance of the latter we choose to use separate CLS tokens in the rest of this study.
@@ 5 Results
In this section, we analyze models obtained with the proposed AM-RADIO framework. First, we touch upon backbone efficiency, then compare with the original teachers (CLIP, DINOv2, SAM), and benchmark models under vision question answering in the LLaVa framework. We will see that the proposed models outperform the original teachers in multiple metrics, including throughput. A sneak peek of the results is presented in Figure 1.
Refer to caption Refer to caption Refer to caption
Benchmarking multiple model architectures in AM-RADIO pipeline ADE20K segmentation example
Figure 2: All models followed the same training protocol. The results from three benchmarks show that RADIO and E-RADIO models outperform others in efficiency. This underperformance in other models might be due to overfitting architectures on supervised ImageNet-1K training. E-RADIO notably delivers results 10 times faster and with a 20% improvement over teacher models. Segmentation results demonstrate that SAM lags behind DINOv2 in segmentation tasks, likely because it is tasked with learning multiple plausible masks, which requires features for object interiors to vary relative to edges.
@@ 5.1 Efficient Students
Table 6: Comparison of backbones. Throughput is measured using TensorRT 9.0.1 on A100 in mixed FP16/FP32 precision at batch size 128 on 224x224 resolution. Rows are sorted by descending throughput order. FD loss refers to the Feature Distillation training loss against the DINOv2 teacher, which exhibits high correlation with the ADE20k mIoU. Bolded models are on the speed/quality Pareto front.
@@
We aim to find an efficient model architecture to speed up the inference of VFM. There are a number of architectural designs aimed at high throughput on GPU devices. Unfortunately, most studies employ varying training recipes, even when training and evaluating on the ImageNet-1K dataset, making comparisons misleading. We use our distillation framework to evaluate several backbones with no change in training hyperparameters.
@@
Upon reviewing the literature on efficient vision backbones focused for high GPU throughput, we pick the following list of architectures: EfficientNetV2[53], ResNetv2[52], RegNetY[47], FasterViT[21], EfficientViT[7], ConvNext[39], NFNet[6], SwinV2[38], MaxViT[54], PoolformerV2[63] and MViTV2[34]. We train all the backbones via distillation on the ImageNet-21k dataset, using OpenCLIP ViT-H/14 (laion2B-s32B-b79K) and DINOv2 G/14 as teachers. We train for 600k steps at a total batch size of 4096, which is empirically deemed enough to reach convergence. Results are compiled in Table 6.
@@
We observe that many models lag behind teachers. Additionally, CNN-like models are significantly faster than ViTs, while the latter are more accurate. The relatively low performance of existing efficient backbones on the dense ADE20k segmentation task is not unexpected since all of them apply a spatial dimension reduction factor of 32 for a final feature maps: the size is 7x7px for input resolution of 224x224px. Thus hardly capable of capturing fine-grain spatial information.
@@
E-RADIO: To overcome this issue, we propose a novel hybrid architecture, named E-RADIO (Efficient RADIO). This design borrows ideas from existing literature and includes an input stem with strided convolutions to downsample the input image by 4x. It then proceeds with 2 stages of YOLOv8 C2f convolution blocks and 2 stages of transformer. For the transformer variant we pick windowed attention (like in SWIN [38]), and interleave local windowed attention with “global” windowed attention as done in [21] and ViTDet [33]. To perform “global” attention we first downsample the feature map by 2x, apply windowed attention, and then upsample the feature maps back to the original resolution. Up-/down-sampling is performed by strided convolution with a kernel size 3x3 and a stride of 2. The last idea is borrowed from EdgeViT [44], which uses local-global-local attention. Details of the architecture are in the Appendix.
Feature map upsample: E-RADIO upsamples final features maps by 2x via a deconvolutional layer and adds them to feature maps from the third stage, resulting in only a 16x spatial resolution reduction. Such upsampling gives an improvement in dense task while being only 10% slower.
Results of E-RADIO in Table 6 demonstrate that the proposed architecture significantly outperforms the competition, and can be seen as an efficient replacement for the much slower full ViT. We pick E-RADIO as our efficient candidate for further study. We envision our framework to be a benchmarking suite for future efficient vision backbone research.
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11 67.31
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
OpenAI CLIP [46] ViT-L/14 1126 75.54 79.80 36.51 67.04 68.80 67.03 53.25 76.47
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-g/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 14 N/A 20.41 26.56 34.34 61.52 63.36 43.29 64.91
RADIO (Ours) E-RADIO-L 3697 77.87 83.73 45.50 79.95 66.22 65.54 43.51 70.74
${}^{\dagger}$ 556 82.73 85.29 50.32 85.17 69.17 67.08 54.92 77.88
Table 7: Comparison of teachers and RADIO models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIOU on the ADE20K [68] and PascalVOC2012 validation datasets. RADIO outperforms all teacher models on most metrics, and then E-RADIO-L is about 6.6x faster than ViT-H/14 while attaining similar metrics on certain tasks.
${}^{\dagger}$e released RADIOv1 model weights after initial publication. We used DFN CLIP [16] (84.37 Zero-Shot), OpenAI CLIP [46], and DINOv2 [43] to train this model. It additionally incorporates 5 register tokens [12]. Refer to table 11 for the original publication numbers.
@@ 5.2 Comparison with teachers
A comprehensive set of results is presented in Table 7. At first, we evaluate teacher models: OpenCLIP, MetaCLIP, and DINOv2. We notice that MetaCLIP is better than OpenCLIP and therefore choose it. DINOv2 provides important properties for dense tasks: ADE20k and VOC. Our E-RADIO-L model is 6.6x faster than MetaCLIP and 11.1x faster than DINOv2-ViT-g/14. At the same time, it outperforms MetaCLIP on k-NN, ADE20k, VOC, and GQA, while also enabling Zero-shot capability that is absent in DINOv2. Our full model, RADIO-ViT-H/14, is as fast as the teachers but outperforms them on 6 out of 8 tasks, demonstrating the efficiency of the proposed distillation framework. Details of these experiments can be found in the next sections.
5.3 Drop-In SAM Replacement
Following [7], we use their evaluation harness to compute the mIOU for instance segmentation using pretrained SAM with vision encoder replaced by our model. Table 8 shows the results of the COCO Instance Segmentation task using the baseline SAM models, and RADIO-ViT.
@@ 5.3 Drop-In SAM Replacement
SAM Base 75.78 29.18
Large 77.01 13.52
Huge 77.18 8.56
RADIO (ours) ${}^{\dagger}$ 75.17 8.20
${}^{\dagger}$ 75.31 8.04
${}^{\dagger}$ 75.42 7.45
Other benchmarks
Family Arch Zero-Shot k-NN ADE20k
SAM ViTDet-L/16 20.41 26.56 34.34
ViTDet-H/16 22.39 28.17 32.16
RADIO (ours) ViTDet-H/14 76.17 83.34 49.01
Table 8: In our SAM replacement experiments using the COCO 2017 instance segmentation framework, we substitute SAM’s vision encoder with our agglomerative (RADIO) model. This model aligns with features just before the encoder’s bottleneck layer. We also examine the impact of varying ViTDet window sizes. Our model almost matches SAM’s results, primarily as a multi-teacher distillation byproduct. Further, when comparing metrics on ImageNet1K and ADE20k, our RADIO model shows superiority.
5.4 Comparison with SAM-CLIP [55]
Concurrently with our work, SAM-CLIP was introduced as a method of fusing SAM and CLIP into a single model. Due to the concurrency of effort, we don’t compare our model with the full suite of metrics demonstrated in their method, however, we do have some overlap in key metrics such as Zero-Shot ImageNet1k, and ADE20k semantic segmentation via linear probing. We present the comparison in table 9, however we note that there are enough differences between these two models that we can’t conclude one way or another what is the superior approach. Instead we’ll argue that DINOv2 does a better job of ADE20k linear probing than SAM, and thus our significantly higher quality on this metric is likely due to the inclusion of DINOv2, which is a key introduction with our approach.
Family Model Zero-Shot ADE20k
SAM ViTDet-H/16 28.2
CLIP ViT-L/14 79.2 36.4
SAM-CLIP ViTDet-B/16 71.7 38.4
RADIO ViT-H/14 78.6 49.0
Table 9: We compare our common key metrics with those demonstrated in SAM-CLIP [55]. We note that there are numerous differences between the two approaches, including model capacity and architecture. SAM-CLIP uses the ViT-B variant of SAM as a starting point, which implies it’s a ViTDet-B/16 architecture. As a result of this choice, their metrics are computed at a resolution of 1024x1024. RADIO trains a vanilla ViT-H/14 from scratch, and as a result of the flexibility gained via the CPE method, we evaluate Zero-Shot ImageNet1k at a resolution of 224x224, and we run ADE20k linear probing at a resolution of 518x518 using the exact same weights. We note that Zero-Shot quality is largely determined by the quality of the CLIP teacher and the capacity of the student. We attribute our superior quality on ADE20k semantic segmentation largely to our inclusion of DINOv2 as a teacher.
Figure 3: VQA examples from LLaVa-1.5 framework with swapped vision encoder. RADIO and E-RADIO are better at capturing fine details, and it is better at counting. We observed that MetaCLIP, DINO and RADIO have high correlation of answers at the level of 0.78, while E-RADIO has 0.68 with others and tends to choose a different answer.

### v4

@@ 3.4.1 Loss Balancing
Due to the number of possible combinations of loss weights between the different teachers, and even which teachers, and possible formulations of loss functions, we mostly opted toward naive loss balancing with all teachers equally weighted for spatial features ($\gamma_{i}=1$). For summary features, we have $\lambda_{CLIP}=\lambda_{DINO}=1$ and $\lambda_{SAM}=0$.
We did experiment with automatic loss balancing using predicted uncertainty [11], AdaLoss [25] (momentum 0.99) and separately with AMTML-KD [39], as ways to learn the balance of $\lambda_{i}$ and $\gamma_{i}$. In the case of AMTML-KD, the model would always collapse its entire weight around the CLIP teacher and would yield worse results than naive manual balancing. Based on the results in table 4, there is very little advantage to the more exotic balancing schemes, so we opt for the "Naive" method throughout the rest of the paper.
@@
Table 4: Loss term balancing methods comparison. We use a ViT-B/14 student, and CLIP+DINOv2 teachers. We found that AdaLoss produces the best results on the ImageNet tasks, but the worst on ADE20K.
@@ 4 Implementation Details
Performing heterogeneous multi-teacher distillation is not trivial due to a mismatch in feature dimensions, input resolutions, concepts for loss computation, and downsampling ratios, as well as challenges in fitting multiple teachers into a single GPU.
General. We train all student models using the AdamW [42] optimizer, batch size 1024, cosine annealing learning rate schedule and base learning rate of $0.001$. We train for 600k steps, resulting in 614M total examples seen. For our best student model, we train using DFN CLIP ViT-H/14 378px, OpenAI CLIP ViT-L/14 336px, DINOv2 ViT-g/14 224px, and SAM ViTDet-H 1024px. We apply random scale + cropping to both student and teacher inputs. We chose the DataComp-1B dataset due to it having the highest quality results of the web-scale datasets we had access to. We train in two stages, first with CLIP+DINOv2 for 300k steps at 256px, and second with CLIP+DINOv2 at 432px plus SAM at 1024px for 300k steps.
@@
  • Standard ViT [15] architecture to match the architecture of teachers. Our best model is a ViT-H/16.
  • Efficient architecture variants prioritizing high throughput on GPUs. See Section 5.1.
Multi-scale Teachers. We choose ViT-H/16 architecture for our student model. To match resolution of SAM features, we feed the expected resolution of $1024^{2}$. Given that our CLIP and DINOv2 teachers are patch-14 models, we opt to feed the student $432^{2}$ inputs, as that is the same effective resolution as $378^{2}$ for patch-14. We found that interpolating DINOv2 features doesn’t degrade results, so the teacher operates at 224px and we upsample the outputs to match the student.
Rank/Teacher Partitioning. We group teacher models by (batch_size, student_resolution), and then distribute the groups to different GPUs, such that each GPU processes a consistent batch size and input resolution. We also sample groups at different rates. For our training setups that include SAM, we train with 64 GPUs, half of which get the CLIP+DINOv2 group with batch size 32 per GPU and input resolution 432, and the other half get SAM with batch size 2 per GPU and input resolution 1024. This results in an effective batch size of 1,152. For CLIP+DINOv2 training, we use 32 GPUs, resulting in batch size 1024.
Multi-Resolution ViTs. Many of our student models use ViT [15] as the base vision architecture. Traditionally, ViTs use a learned position embedding for each input patch in an image, which in turn enforces that the model always operates at a constant resolution. We employ the Cropped Position Embedding (CPE) [30] augmentation with the number of positions being equal to $128^{2}$. The position embeddings are then randomly cropped and interpolated to match the number of input patches for the student model. Even when training with CLIP+DINOv2 at 224 resolution, we found that this technique results in a negligible drop (Table 5) in summary metrics, but improved semantic segmentation linear probing mIOU. For heterogeneous-resolution students, this is a seamless technique that allows ViT to operate at arbitrary resolutions within some envelope.
@@
Non-CPE 82.96 47.30
CPE 82.84 48.52
Table 5: Comparing identical ViT-L/14 student models, with and without CPE [30] formulation. While the student only ever trains at $224^{2}$ resolution, CPE allows us to generalize to $518^{2}$ resolution, not only improving over non-CPE, but even outperforming DINOv2-g itself.
High-Resolution ViT Student. In SAM, they employ the ViTDet [34] architecture as a way to reduce the computational and memory burden of ViT models at high-resolution. We reformulate this arch instead into a training augmentation, where we sample a window size from a set of possible window sizes. This allows us to reduce the computational burden of training the student model with the SAM teacher, and, as we make the window size flexible, it provides an additional throughput scaling mechanism during inference. Table 8 demonstrates our ability to replace SAM’s encoder. Separately, we found that high resolution training was unstable, so we apply spectral reparametrization [68] and a weight decay of $0.02$ to prevent attention entropy collapse.
Student/Teacher Resolution Mismatch. When the student and teacher downsample images through their processing stack at different rates, it results in the output feature vectors having different resolutions. For example, if the teachers use a ViT-H/14 architecture and student a ViT-H/16, it means that the student outputs a $14^{2}$ feature map, and the teachers a $16^{2}$ feature map. For $L_{\text{features}}$ we bilinearly interpolate the outputs to match the larger resolution between the student and teacher features.
Zero Shot k-NN ADE20K VOC VQAv2
CLS token 78.55 83.91 49.01 83.51 77.66
Avgpool 80.12 83.83 38.36 77.04 78.28
Table 6: Comparing identical ViT models, with CLS token and average pooling summarization.
Feature Summarization. In 3.4 we explained how teacher summary features are extracted using the “class token” of their respective ViT models. We now turn our attention to the summarization of student features. ViTs have 2 options: (i) a separate summarization “CLS” token or (ii) average pooling patch tokens. We evaluate both options in Table 6. We observe that average pooling improves summary loss, but has a more significant detrimental effect on the feature loss. Given the importance of the latter we choose to use separate CLS tokens.
@@ 5 Results
In this section, we analyze models obtained with the proposed AM-RADIO framework. First, we touch upon backbone efficiency, then compare with the original teachers (CLIP, DINOv2, SAM), and benchmark models under vision question answering in the LLaVa framework. We will see that the proposed models outperform the original teachers in multiple metrics, including throughput. Results are shown in Figure 2 and Table 1.
Figure 3: All models followed the same training protocol. The results from three benchmarks show that RADIO and E-RADIO models outperform others in efficiency. This under-performance in other models might be due to overfitting architectures on supervised ImageNet-1K training. E-RADIO notably delivers results 10 times faster and with a 20% improvement over teacher models. We study E-RADIO at 224px resolution, with a window size of 7.
@@ 5.1 Efficient Students
Table 7: Comparison of backbones. Throughput is measured using TensorRT 9.0.1 on A100 in mixed FP16/FP32 precision at batch size 128 on $224^{2}$px resolution. Sorted by descending throughput order. FD loss is the Feature Distillation training loss against the DINOv2 teacher, it exhibits high correlation with the ADE20k mIoU. Bolded models form the speed/quality Pareto front.
@@
We aim to find an efficient model architecture to speed up the inference of VFM. There are a number of architectural designs aimed at high throughput on GPU devices. We use our distillation framework to evaluate several backbones with no change in training hyperparameters.
@@
Upon reviewing the literature on efficient vision backbones focused for high GPU throughput, we pick the following list of architectures: EfficientNetV2 [55], ResNetv2 [54], RegNetY [49], FasterViT [22], EfficientViT [7], ConvNext [41], NFNet [6], SwinV2 [40], MaxViT [56], PoolformerV2 [65] and MViTV2 [35]. We train all the backbones via distillation on the ImageNet-21k dataset, using OpenCLIP ViT-H/14 (laion2B-s32B-b79K) and DINOv2 g/14 as teachers. Results are compiled in Table 7.
@@
We observe that many models lag behind teachers. Additionally, CNN-like models are significantly faster than ViTs, while the latter are more accurate. The relatively low performance of existing efficient backbones on the dense ADE20k segmentation task is not unexpected since all of them apply a spatial dimension reduction factor of 32 for final feature maps of size $7^{2}$ for input resolution of $224^{2}$px, thus hardly capable of capturing fine-grain spatial information.
@@
E-RADIO: To overcome this issue, we propose a novel hybrid architecture, named E-RADIO (Efficient RADIO). This design borrows ideas from existing literature and includes an input stem with strided convolutions to downsample the input image by 4x. It then proceeds with 2 stages of YOLOv8 C2f convolution blocks and 2 stages of transformer. For the transformer variant we pick windowed attention (like in SWIN [40]), and interleave local windowed attention with “global” windowed attention as done in [22] and ViTDet [34]. To perform “global” attention we first downsample the feature map by 2x, apply windowed attention, and then upsample the feature maps back to the original resolution. Up-/down-sampling is performed by strided convolution with a kernel size 3x3 and a stride of 2. The last idea is borrowed from EdgeViT [46], which uses local-global-local attention. See Appendix for details. Finally, E-RADIO upsamples final feature maps by 2x via a deconvolutional layer and adds them to feature maps from the third stage, resulting in only a 16x spatial resolution reduction. Such upsampling gives an improvement in dense task while being only 10% slower. Results of E-RADIO in Table 7 demonstrate that the proposed architecture significantly outperforms the competition, and can be seen as an efficient replacement for the much slower full ViT.
@@ 5.2 Comparison with teachers
A comprehensive set of results is presented in Table 1. We notice that MetaCLIP is better than OpenCLIP, and DFN CLIP better than MetaCLIP. DINOv2 provides important properties for dense tasks: ADE20k and VOC. Our E-RADIO-L model is significantly faster than all ViT models. At the same time, it strongly outperforms MetaCLIP on most metrics at matched throughput, while also enabling Zero-shot capability that is absent in DINOv2 and SAM. Our full model, ViT-H/16, is as fast as the teachers but outperforms them on 6 out of 9 tasks, demonstrating the efficiency of the proposed distillation framework.
Drop-In SAM Replacement. Following [7], we use their evaluation harness to compute the mIOU for instance segmentation using pretrained SAM with vision encoder replaced by our model. Table 8 shows the results of the COCO Instance Segmentation task using the baseline SAM models and RADIO.
@@
SAM Base 75.78 50.94
Large 77.02 20.62
Huge 77.18 11.83
E-RADIO (ours) Large 76.31 121.74
RADIO (ours) ViTDet-H/16-W8 76.09 29.09
ViTDet-H/16-W16 76.23 27.91
Table 8: We substitute SAM’s vision encoder with our RADIO model. RADIO aligns with SAM’s features just before the encoder’s neck layer. We also examine the impact of varying ViTDet window sizes. Differences in throughput owe to the fact that RADIO doesn’t use relative positional embeddings and we reduced shuffling with our patch reordering algorithm (in appendix). Throughput is computed on an NVIDIA A100 GPU using TensorRT and batch size 16. This is the same model, just with a different window size setting.